zarkam21 (zarkam21):

65=1/2(adc-130) What is adc?

OpenStudy (zyberg):

Multiply both sides by 2, then add 130 to both sides and you will have your answer ;)

zarkam21 (zarkam21):

By 2 or 1/2

OpenStudy (zyberg):

By 2

OpenStudy (zyberg):

That was my mistake back then ;) In other words you are dividing both sides by 1/2.

zarkam21 (zarkam21):

Well for 65, it would be 32.5

OpenStudy (zyberg):

65*2 = 130 65/(1/2)=130

zarkam21 (zarkam21):

Okay so the other side would be 130 right, because 130*2=260 and 260/1/2=130

jhonyy9 (jhonyy9):

1 65 = --- * (adc -130) 2 this is your equation right wrote ?

zarkam21 (zarkam21):

Yes

jhonyy9 (jhonyy9):

so from this result 2*65 = adc -130 yes ?

OpenStudy (zyberg):

65=1/2(adc-130) 130 = adc - 130 adc = 260

OpenStudy (zyberg):

65/ (1/2) = ((1/2)/(1/2))*(adc -130) 130 = adc - 130 260 = adc - 130 +130

zarkam21 (zarkam21):

Ohh okay

jhonyy9 (jhonyy9):

@Zyberg please dont give directly right answers please - this is the rule of OS @mathmate please

zarkam21 (zarkam21):

No @Zyberg did not give a direct answer, its help

OpenStudy (zyberg):

@jhonyy9 I didn't. I told her how to get the answer and she said that it's 260, but she obtained it in a wrong way, so I showed her the right way.

zarkam21 (zarkam21):

She showed me the process

jhonyy9 (jhonyy9):

ok. sorry

zarkam21 (zarkam21):

So this would be 207

zarkam21 (zarkam21):

27=1/2(dgf-153)

zarkam21 (zarkam21):

@Zyberg

zarkam21 (zarkam21):

@jhonyy9

jhonyy9 (jhonyy9):

what is the question here ?

zarkam21 (zarkam21):

27=1/2(dgf-153)

zarkam21 (zarkam21):

I tried the problem out and got the answer 207, is this correct??

zarkam21 (zarkam21):

@Zyberg is 207 correct for 27=1/2(dgf-153)

OpenStudy (zyberg):

@zarkam21 Yes! :)